Sonya inherited $ 35,000. She invested some of the money at 2.5 % simple interest and the rest of 3.5 % simple interest. In the first year, she earned
$ 1,075 interest. How many dollars did she earn from her investment at 3.5 % interest ?

Answer by checkley79(3341)

.035X+.025(35,000-X)=1.075
.035X+875-.025X=1.075
.010X=1,075-875
.010X=200
X=200/.010
X=20,000 AMOUNT INVESTED @ 3.5%.
35,000-20,000=15,000 AMOUNT INVESTED @ 2.5%.PROOF:
.035*20,000+.025*15,000=1,075
700+375=1,075
1,075=1,075